OpenStudy (anonymous):

What is the value of the coefficient "b" when the quadratic equation y = (3x − 5)(2x + 3) is written in standard form? 19 6 −15 −1

hero (hero):

Hint: Multiply (3x-5)(2x+3)

OpenStudy (anonymous):

i did but i got b=-9

hero (hero):

Well double check your work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got the same answer

hero (hero):

y= (3x-5)(2x+3) = 2x(3x-5) + 3(3x-5) = 2x^2 - 10x + 9x - 15 = 2x^2 - x - 15